OpenStudy (anonymous):

A vase contains 12 tulips and 20 roses. What is the ratio of roses to the total number of flowers in the vase? A. 3 to 8 B. 3 to 5 C. 5 to 8 D. 5 to 3

OpenStudy (anonymous):

@AlexandervonHumboldt2 @AloneS @GretaKnows @Jaynator495 @just_one_last_goodbye @mathstudent55 @nincompoop @pooja195 @Purple_Snack @purple_pink

OpenStudy (campbell_st):

start by using the given numbers in the correct order tulips Roses 12 20 they are both even number so halve each... if the answer is still 2 even numbers, halve them again...

OpenStudy (purple_pink):

i think b, but wait

OpenStudy (anonymous):

ok

OpenStudy (purple_pink):

i need to see if im right, give me 1minute

OpenStudy (anonymous):

ok

OpenStudy (purple_pink):

yes it is B

OpenStudy (mathstudent55):

You need to first see what the question is. You are looking for a ratio. A ratio can be written as a fraction. \(\dfrac{~~~~~~~~~~}{}\) Above is a fraction line. Now we see what to write on the fraction.

OpenStudy (mathstudent55):

The question is: Ratio of roses to total number of flowers. We know the number of roses because we are given that number. \(\dfrac{number~of~roses}{total~number~of~flowers } = \dfrac{20} {}\) We need to total number of flowers. 12 tulips + 20 roses = 32 flowers. \(\dfrac{number~of~roses}{total~number~of~flowers } = \dfrac{20} {32}\)

OpenStudy (mathstudent55):

Now we reduce 20/32 until we can't reduce it anymore. \(\dfrac{number~of~roses}{total~number~of~flowers } = \dfrac{20} {32} = \dfrac{10}{16} = \dfrac{5}{8}\)

OpenStudy (anonymous):

thanks gu]ys

OpenStudy (mathstudent55):

Notice our answers are different. You need to pick the correct one.

OpenStudy (mathstudent55):

Answer is not B.

OpenStudy (mathstudent55):

The question is NOT the ratio of tulips to roses. The question is the ratio of roses to total flowers.